OpenStudy (anonymous):

Algebra 1: please please help, I only have 5 minutes left and I have no idea how to get the answer A system of linear inequalities is shown below. x - y > 3 y + x ≤ 2 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (shiraz14):

1. Draw the graph of x-y = 3 (i.e. the line y=x-3). Since x-y > 3, you will shade the region on the graph above this line. 2. Draw the graph of y+x = 2 (i.e. the line y=-x+2). Since y+x ≤ 2, you will shade the region on the graph below this line (and include this line). Find their overlapping regions, and you should find the solution set to this question.

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (shiraz14):

You're welcome.